the perimeter of a rectangle is 68cm. the length is 4 cm than the width identify your two variables. formulate two equations.
Answer by rajagopalan(174)

Let the Lenth =L
Let the width=B
Equation 1 ...L=B+4
Equation 2 ...2(L+B)=68
*****
For solving
substitute L=B+4 in eqn 2
You get 2(B+4+B)=68
2(2B+4)=68
2B+4=68/2
2B=34-4=30
B=15
L=19